OpenStudy (anonymous):

4 / a-3 = 2/5 Solve please

OpenStudy (anonymous):

a = 13

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

4*5/2 = a -3 => 10 = a -3 => a = 10+3 =13

OpenStudy (anonymous):

did I solve for the proportion there?

OpenStudy (anonymous):

yes..........

OpenStudy (anonymous):

Iqht thanks

OpenStudy (anonymous):

u welcome